# Steady state diagram of interacting fermionic atoms coupled to dissipative cavities

### v1

@@ Abstract
We investigate fermionic atoms subjected to an optical lattice and coupled to a high finesse optical cavity with photon losses. A transverse pump beam introduces a coupling between the atoms and the cavity field. We explore the steady state phase diagram taking fluctuations around the mean-field of the atoms-cavity coupling into account. Our approach allows us to investigate both one- and higher-dimensional atomic systems. The fluctuations beyond mean-field lead to an effective temperature which changes the nature of the self-organization transition. We find a strong dependence of the results on the atomic filling, in particular when contrasting the behavior at low filling and at half filling. At low filling the transition to a self-organized phase takes place at a critical value of the pump strength. In the self-organized phase the cavity field takes a finite expectation value and the atoms show a modulation in the density. Surprisingly, at even larger pump strengths a strongly non-monotonous behavior of the temperature is found and hints towards effects of cavity cooling at many-body resonances. Additionally multiple self-organized stable solutions of the cavity field and the atoms occur, signaling the presence of a fluctuation-induced bistability, with the two solutions having different effective temperatures previously discussed in Ref. Tolle et al. (2025a). In contrast, at half filling a bistable region arises at the self-organization transition already neglecting the fluctuations. The presence of the fluctuations induce an effective temperature as at lower filling and change the behavior of the transition and the steady states drastically. We analyze the properties of the occurring steady states of the coupled atoms-cavity system.
@@ I Introduction
The recent experimental achievements of ultracold atoms in optical cavities allow one to access the strong coupling regime and gain precise control over the collective matter-light coupling Ritsch et al. (2013); Mivehvar et al. (2021). Bosonic atoms have been experimentally coupled to cavities, either as three-dimensional Bose-Einstein condensates Baumann et al. (2010); Kroeze et al. (2018); Vaidya et al. (2018); Keßler et al. (2021); Ferri et al. (2021); Dreon et al. (2022); Kongkhambut et al. (2022), or confined to an external optical lattice structure Landig et al. (2016); Hruby et al. (2018); Klinder et al. (2015a, b). In their fermionic counterpart both the steady state self-organization phase transition to density wave states Helson et al. (2023); Zwettler et al. (2025a); Zhang et al. (2021) and the out-of-equilibrium dynamics have been experimentally investigated Zwettler et al. (2025b); Wu et al. (2023). Coupling fermionic atoms to optical cavities has been a rich playground also theoretically Larson et al. (2008a); Sun et al. (2011); Piazza and Strack (2014a); Chen et al. (2014); Keeling et al. (2014); Chen et al. (2015); Wolff et al. (2016); Marijanović et al. (2024); Tolle et al. (2025a); Ortuño-Gonzalez et al. (2025). In particular, the dissipative dynamics has been proposed for realization of exotic quantum phases Dong et al. (2014); Fan et al. (2018); Colella et al. (2018, 2019); Schlawin et al. (2019); Zheng and Wang (2020); Nie and Zheng (2023); Mivehvar et al. (2017), topological effects Kollath et al. (2016); Sheikhan et al. (2016a, b); Mendez-Cordoba et al. (2020); Pan et al. (2015), and pairing in superfluids Schlawin and Jaksch (2019). While a large part of the theoretical atoms-cavities studies have relied on a mean-field treatment of the cavity field Ritsch et al. (2013); Mivehvar et al. (2021), a growing body of the literature is showing the importance of including the quantum fluctuations of the atoms-cavity coupling term in order to correctly capture the steady states and the dissipative dynamics Gopalakrishnan et al. (2010); Piazza and Strack (2014b); Wall et al. (2016); Damanet et al. (2019); Lin et al. (2019); Halati et al. (2020a, b); Bezvershenko et al. (2021); Halati et al. (2022); Link et al. (2022); Chiriacò et al. (2022); Jäger et al. (2022); Lentrodt et al. (2023); Halati et al. (2025a); Tolle et al. (2025a); Halati et al. (2025b); Müller and Strunz (2025); Schmit et al. (2025); Orso et al. (2025).
@@
In this work, we compute the steady state phase diagram for interacting fermionic atoms confined to optical lattices and transversely coupled to the field of a dissipative optical cavity over a wide range of parameters. We employ an approach based on the many-body adiabatic elimination method that takes into account the quantum fluctuations which was originally developed for bosonic atoms Bezvershenko et al. (2021). This approach goes beyond the mean-field treatment of the cavity field. We compare the findings of the two approaches and point out similarities and differences in the properties of the steady state diagram. Our work here complements the investigation of Ref. Tolle et al. (2025a) in which we focused on the existence of a fluctuation-induced bistability. Additionally to the previous work, we analyze in detail the self-organization transition and the cavity induced cooling due to many-body resonances. We derive analytical formulas for the critical coupling in different regimes. We summarize also the findings of the fluctuation induced bistability at low filling of Ref. Tolle et al. (2025a) for completeness and give additional details. Going beyond the previous work, we analyze the self-organization transition at half filling, which shows a bistability which is already present in the zero-temperature mean-field approach and thus, has a very different character than the fluctuation-induced bistability present at low filling.
@@
The paper is organized as follows. We begin by introducing our model of interacting fermions on an external optical lattice that globally couples to a dissipative cavity mode via transverse driving with a pump laser (Sec. II). Then we proceed to introduce the analytical methods, we derive to determine the steady state phase diagram (Sec. III). We employ the many-body adiabatic elimination technique that takes into account perturbatively the quantum fluctuations of the atoms-cavity coupling (Sec. III.2). The fluctuations determine a self-consistent effective temperature of the atoms. We provide a stability condition for the thermal steady state solutions (Sec. III.3). Further, we derive a low-hopping thermodynamic limit that can provide results in higher dimensions (Sec. III.4). In order to facilitate the understanding of the phases we see in the complex atoms-cavity coupled system, we provide an overview of the ionic Hubbard model (Sec. IV), since the effective Hamiltonian in our mean-field approach takes this form with a self-consistently determined staggered potential. We then proceed to present our results for the steady state phase diagrams across the self-ordering transition at both low filling (Sec. V.1) and commensurate half filling (Sec. V.2). Special focus is put on cooling processes. We extend the concept of cavity-cooling (Sec. V.1.2) to the many-body regime (Sec. V.1.3) and show that an efficient transfer of energy between the photons and the atoms, mediated by the fluctuations, determines many of the features present in the phase diagram, including the fluctuation-induced bistability (Sec. V.1.4) we introduced in Ref. Tolle et al. (2025a). Here we contrast the fundamentally different nature of the bistability at low fillings and the special case of commensurate fillings (Sec. V.2.1).
@@ II Setup and Model
The density matrix $\hat{\rho}$ contains both the photonic and atomic degrees of freedom. The cavity experiences photon losses described by the dissipator with amplitude $\varGamma$ and the photon annihilation jump operator $\hat{a}$. The atoms-cavity system is described by the following Hamiltonian Ritsch et al. (2013); Mivehvar et al. (2021); Tolle et al. (2025a)
@@ III Methods
In this section, we begin by introducing the mean-field elimination of the cavity mode, reducing the system to an effective atomic model (Sec. III.1). Then we go beyond this approximation by including fluctuations in the atoms-cavity coupling with our many-body adiabatic elimination approach (Sec. III.2). Further, a stability condition for the solutions from the previously obtained equations under variations of the steady state cavity field is derived (Sec. III.3). We conclude the methods section by calculating a perturbative model in the weak tunneling regime (Sec. III.4), going to the thermodynamic limit.
@@ III.1 Mean-field elimination of the cavity mode
For our model, Eq. (2), the effective Hamiltonian is given by the ionic Hubbard model, whose properties we discuss in Sec. IV. The value of the staggered potential is $\hbar g\lambda$. A common approach is to take the self-consistently determined ground state of the effective Hamiltonian $\hat{H}_{\text{eff}}$ Ritsch et al. (2013); Mivehvar et al. (2021), which in the following we dub the zero-temperature mean-field method (T=0T\!=\!0 MF). However, this is an arbitrary choice since any eigenstate of $\hat{H}_{\text{eff}}$ is a steady state solution of the atomic equations of motion Bezvershenko et al. (2021); Tolle et al. (2025a); Halati (2021).
@@ III.2 Many-body adiabatic elimination technique for atoms-cavity coupled systems
In recent years, the crucial role of fluctuations in the coupling of quantum matter and quantum light has been the focus of several theoretical works Mivehvar et al. (2021); Gopalakrishnan et al. (2010); Piazza and Strack (2014b); Wall et al. (2016); Damanet et al. (2019); Halati et al. (2020a, b); Bezvershenko et al. (2021); Halati et al. (2022); Chiriacò et al. (2022); Jäger et al. (2022); Link et al. (2022); Lentrodt et al. (2023); Halati et al. (2025a); Tolle et al. (2025a); Müller and Strunz (2025); Halati et al. (2025b); Schmit et al. (2025); Orso et al. (2025). Here, we employ a method developed in Ref. Bezvershenko et al. (2021) in the framework of the many-body adiabatic elimination technique García-Ripoll et al. (2009); Reiter and Sørensen (2012); Kessler (2012); Poletti et al. (2013); Zanardi and Campos Venuti (2014); Sciolla et al. (2015); Lange et al. (2018). This approach includes perturbatively the fluctuations in the atoms-cavity coupling on top of the mean-field treatment (see Sec. III.1). This method was developed and benchmarked against numerically exact results for bosonic atoms Bezvershenko et al. (2021); Halati et al. (2022); Halati (2021) and extended to fermionic atoms in Ref. Tolle et al. (2025a).
@@
with $\beta\!=\!1/k_{B}T$ the inverse temperature and $Z$ the partition function. In our case, this is a justified assumption for a finite occupation of the cavity field, since the corresponding effective Hamiltonian $\hat{H}_{\text{eff}}$, Eq. (6), is the ionic Hubbard model. This model was shown to obey Gaussian orthogonal ensemble (GOE) statistics in the presence of the staggered potential De Marco et al. (2022).
@@
is evaluated for the thermal state $\hat{\rho}_{\text{at}}$. The self-consistently determined temperature can stabilize to different values as we vary the parameters throughout the phase diagram, as we discuss in Sec. V. In particular this effect is a crucial ingredient for the many-body-cooling mechanisms and the fluctuation-induced bistability Tolle et al. (2025a). In the normal phase fluctuations in the atoms-cavity coupling are present despite the vanishing steady state cavity field. Nevertheless, one should be careful in one dimension using the thermal state Ansatz given in Eq. (14), as the Fermi-Hubbard model is integrable, meaning that one should use a generalized Gibbs ensemble state including the conservation laws below the transition Halati et al. (2022); Lange et al. (2017, 2018).
@@ III.2.1 Large temperature limit
This result agrees with the result of inverse effective temperature, $\beta$, for bosons Bezvershenko et al. (2021) and is consistent with results based on the semiclassical limit for non-interacting atoms Asbóth et al. (2005); Schütz and Morigi (2014); Piazza and Strack (2014b).
@@ IV.1 Ionic Hubbard Hamiltonian
Here the expectation values are evaluated for the state one is interested in, so for example the ground state, or the finite temperature state. For the ground state one can use $\langle\hat{\gamma}^{\dagger}_{\pm,k,\sigma}\hat{\gamma}_{\pm,k,\sigma}\rangle\!=\!1$ for the filled momenta and $\langle\hat{\gamma}^{\dagger}_{\pm,k,\sigma}\hat{\gamma}_{\pm,k,\sigma}\rangle\!=\!0$ otherwise.
@@ IV.2 Ground state phase diagram and excitations
We discuss the phase diagram and low lying excitations of the ionic Hubbard model in one dimension. Since distinct phases arise at different fillings, we consider here two different fillings $n\!=\!N/L\!=\!\{1/2;1\}$.
@@ V Results
In this section, we present our results on the steady state phase diagram of the atoms-cavity coupled system. The transversely pumped coupled atoms-cavity open system exhibits a self-ordering transition Ritsch et al. (2013); Mivehvar et al. (2021) which has been studied widely for bosonic and fermionic atoms Asbóth et al. (2005); Larson et al. (2008b, a); Niedenzu et al. (2011); Piazza and Strack (2014b); Keeling et al. (2014); Chen et al. (2014). Our results investigate the self-organization transition for the considered system of interacting fermions coupled to the optical cavity. We analyze in detail the case of quarter filling (Sec. V.1), complementing our work presented in Ref. Tolle et al. (2025a). Whereas the shown numerical results are obtained at n=1/2n\!=\!1/2, our analytical findings can be generalized to lower fillings $n\!\leq\!1/2$. Additionally, we present and analyze results at half filling n=1n\!=\!1 (Sec. V.2). Fundamentally different behavior is expected for both fillings.
We begin the discussion (see Sec. V.1) by focusing on the regime around the self-ordering transition at $g\!\sim\!g_{\text{cr}}$ at quarter filling. We contrast the T=0T\!=\!0 MF results to the approach which includes fluctuations in the atoms-cavity coupling (Sec. V.1.1). Analytical approximations for the phase transition are derived in several limits. Then we emphasize features originating from the different cavity-cooling (Sec. V.1.2) and many-body-cooling mechanisms (Sec. V.1.3), that are only possible due to the presence of fluctuations. At quarter filling, this includes the fluctuations-induced bistability (Sec. V.1.4) which has been the focus of Ref. Tolle et al. (2025a). In Sec. V.2.1, we turn to the bistability occurring at half filling. In this case the self-organization transition is coupled with the onset of a bistability which is already visible for the T=0T\!=\!0 MF results.
We compare our results from solving of Eq. (5) and Eq. (15) with exact diagonalization (ED) for finite systems in one dimension to those obtained in the thermodynamic weak-tunneling perturbation Eqs. (29). The latter are also applicable for higher dimensions of the atomic lattice.

### v2

@@ Abstract
We investigate fermionic atoms subjected to an optical lattice and coupled to a high finesse optical cavity with photon losses. A transverse pump beam introduces a coupling between the atoms and the cavity field. We explore the steady state phase diagram taking fluctuations around the mean-field of the atoms-cavity coupling into account. Our approach allows us to investigate both one- and higher-dimensional atomic systems. The fluctuations beyond mean-field lead to an effective temperature which changes the nature of the self-organization transition. We find a strong dependence of the results on the atomic filling, in particular when contrasting the behavior at low filling and at half filling. At low filling the transition to a self-organized phase takes place at a critical value of the pump strength. In the self-organized phase the cavity field takes a finite expectation value and the atoms show a modulation in the density. Surprisingly, at even larger pump strengths a strongly non-monotonous behavior of the temperature is found and hints towards effects of cavity cooling at many-body resonances. Additionally multiple self-organized stable solutions of the cavity field and the atoms occur, signaling the presence of a fluctuation-induced bistability, with the two solutions having different effective temperatures previously discussed in Ref. Tolle et al. (2025a). In contrast, at half filling a bistable region arises at the self-organization transition already neglecting the fluctuations. The presence of the fluctuations induce an effective temperature, as at lower fillings, and change the behavior of the transition and the steady states drastically. We analyze the properties of the occurring steady states of the coupled atoms-cavity system.
@@ I Introduction
The recent experimental achievements of ultracold atoms in optical cavities allow one to access the strong coupling regime and gain precise control over the collective matter-light coupling Ritsch et al. (2013); Mivehvar et al. (2021). Bosonic atoms have been experimentally coupled to cavities, either as three-dimensional Bose-Einstein condensates Baumann et al. (2010); Kroeze et al. (2018); Vaidya et al. (2018); Keßler et al. (2021); Ferri et al. (2021); Dreon et al. (2022); Kongkhambut et al. (2022), or confined to an external optical lattice structure Landig et al. (2016); Hruby et al. (2018); Klinder et al. (2015a, b). In their fermionic counterpart both the steady state self-organization phase transition to density wave states Helson et al. (2023); Zwettler et al. (2025a); Zhang et al. (2021) and the out-of-equilibrium dynamics have been experimentally investigated Zwettler et al. (2025b); Wu et al. (2023). Coupling fermionic atoms to optical cavities has been a rich playground also theoretically Larson et al. (2008a); Sun et al. (2011); Piazza and Strack (2014a); Chen et al. (2014); Keeling et al. (2014); Chen et al. (2015); Wolff et al. (2016); Marijanović et al. (2024); Tolle et al. (2025a); Ortuño-Gonzalez et al. (2025). In particular, the dissipative dynamics has been proposed for realization of exotic quantum phases Dong et al. (2014); Fan et al. (2018); Colella et al. (2018, 2019); Schlawin et al. (2019); Zheng and Wang (2020); Nie and Zheng (2023); Mivehvar et al. (2017), topological effects Kollath et al. (2016); Sheikhan et al. (2016a, b); Mendez-Cordoba et al. (2020); Pan et al. (2015), and pairing in superfluids Schlawin and Jaksch (2019). While a large part of the theoretical atoms-cavities studies have relied on a mean-field treatment of the cavity field Ritsch et al. (2013); Mivehvar et al. (2021), a growing body of the literature is showing the importance of including the quantum fluctuations of the atoms-cavity coupling term in order to correctly capture the steady states and the dissipative dynamics Gopalakrishnan et al. (2010); Piazza and Strack (2014b); Wall et al. (2016); Damanet et al. (2019); Lin et al. (2019); Halati et al. (2020a, b); Bezvershenko et al. (2021); Halati et al. (2022); Link et al. (2022); Chiriacò et al. (2022); Jäger et al. (2022); Lentrodt et al. (2023); Halati et al. (2025a); Tolle et al. (2025a); Halati et al. (2025a); Müller and Strunz (2025); Schmit et al. (2025); Orso et al. (2025).
@@
In this work, we compute the steady state phase diagram for interacting fermionic atoms confined to optical lattices and transversely coupled to the field of a dissipative optical cavity over a wide range of parameters. In particular, we focus on the case of low fillings and on the commensurate half filling case. We employ a commonly used mean-field approach and additionally, the many-body adiabatic elimination method that takes into account the quantum fluctuations of the atoms-cavity coupling. This approach was originally developed for bosonic atoms Bezvershenko et al. (2021) in order to capture the mixed state character of the atomic steady states naturally occurring in the presence of dissipation, when the cavity field fluctuations are taken into account. We compare the findings of the two approaches and point out similarities and differences in the properties of the steady state diagram. Our work here complements the investigation of Ref. Tolle et al. (2025a) in which we focused on the existence of a fluctuation-induced bistability at low filling. We briefly summarize the previous findings of the fluctuation induced bistability at low filling of Ref. Tolle et al. (2025a) for completeness in Sec. V.1.4. Additionally to the previous work, we analyze in detail the self-organization transition and the cavity induced cooling due to many-body resonances. We derive analytical formulas for the critical coupling in different regimes. Further, we analyze the self-organization transition at half filling a case not considered in the previous work. For the commensurate half-filling case the transition exhibits a bistability analog to the optical bistability Ritsch et al. (2013), which has a very different character than the fluctuation-induced bistability present at low filling. We uncover that the bistability at half filling is already present in the zero-temperature mean-field approach. However, we show that important deviations occur when fluctuations are included, in particular, when the cavity energy scales dominate.
@@
The paper is organized as follows. We begin by introducing our model of interacting fermions on an external optical lattice that globally couples to a dissipative cavity mode via transverse driving with a pump laser (Sec. II). Then we proceed to introduce the analytical methods, we derive to determine the steady state phase diagram (Sec. III). We employ the many-body adiabatic elimination technique that takes into account perturbatively the quantum fluctuations of the atoms-cavity coupling (Sec. III.2). The fluctuations determine a self-consistent effective temperature of the atoms. We provide a stability condition for the thermal steady state solutions (Sec. III.3). Further, we obtain perturbative results for vanishing kinetic energy in the thermodynamic limit, which is also applicable in higher dimensions (Sec. III.4). In order to facilitate the understanding of the phases we obtain in the complex atoms-cavity coupled system, we provide an overview of the ionic Hubbard model (Sec. IV), since the effective Hamiltonian in our mean-field approach takes this form with a self-consistently determined staggered potential. We then proceed to present our results for the steady state phase diagrams across the self-ordering transition at both low filling (Sec. V.1) and commensurate half filling (Sec. V.2). Special focus is put on cooling processes. We extend the concept of cavity-cooling (Sec. V.1.2) to the many-body regime (Sec. V.1.3), and show that an efficient transfer of energy between the photons and the atoms, mediated by the fluctuations, determines many of the features present in the phase diagram, including the fluctuation-induced bistability (Sec. V.1.4) we introduced in Ref. Tolle et al. (2025a). Here we contrast the fundamentally different nature of the bistability at low fillings and the special case of commensurate fillings (Sec. V.2.1).
@@ II Setup and Model
The density matrix $\hat{\rho}$ contains both the photonic and atomic degrees of freedom. The cavity experiences photon losses described by the dissipator with the amplitude $\varGamma$ and the photon annihilation jump operator $\hat{a}$. The atoms-cavity system is described by the following Hamiltonian Ritsch et al. (2013); Mivehvar et al. (2021); Tolle et al. (2025a)
@@ III Methods
In this section, we begin by introducing the mean-field elimination of the cavity mode, reducing the system to an effective atomic model (Sec. III.1). Then we go beyond this approximation by including fluctuations in the atoms-cavity coupling with our many-body adiabatic elimination approach (Sec. III.2). Further, a stability condition for the solutions from the previously obtained equations under variations of the steady state cavity field is derived (Sec. III.3). We conclude the methods section by performing a perturbative calculation in the weak tunneling regime (Sec. III.4), going to the thermodynamic limit.
@@ III.1 Mean-field elimination of the cavity mode
For our model, Eq. (2), the effective Hamiltonian is given by the ionic Hubbard model, whose properties we discuss in Sec. IV. The value of the staggered potential is $\hbar g\lambda$. A common approach is to take the self-consistently determined ground state of the effective Hamiltonian $\hat{H}_{\text{eff}}$ Ritsch et al. (2013); Mivehvar et al. (2021), which in the following we dub the zero-temperature mean-field method (T=0T\!=\!0 MF). However, this is an arbitrary choice since any eigenstate of $\hat{H}_{\text{eff}}$ is a steady state solution of the atomic equations of motion Bezvershenko et al. (2021); Tolle et al. (2025a); Halati (2021). Thus, in order to determine the contributions of each eigenstate of $\hat{H}_{\text{eff}}$ to the steady state, one needs to go beyond the mean-field approach.
@@ III.2 Many-body adiabatic elimination technique for atoms-cavity coupled systems
In recent years, the crucial role of fluctuations in the coupling of quantum matter and quantum light has been the focus of several theoretical works Mivehvar et al. (2021); Gopalakrishnan et al. (2010); Piazza and Strack (2014b); Wall et al. (2016); Damanet et al. (2019); Halati et al. (2020a, b); Bezvershenko et al. (2021); Halati et al. (2022); Chiriacò et al. (2022); Jäger et al. (2022); Link et al. (2022); Lentrodt et al. (2023); Halati et al. (2025a); Tolle et al. (2025a); Müller and Strunz (2025); Halati et al. (2025a); Schmit et al. (2025); Orso et al. (2025). Here, we employ a method developed in Ref. Bezvershenko et al. (2021) in the framework of the many-body adiabatic elimination technique García-Ripoll et al. (2009); Reiter and Sørensen (2012); Kessler (2012); Poletti et al. (2013); Zanardi and Campos Venuti (2014); Sciolla et al. (2015); Lange et al. (2018). This approach perturbatively includes the fluctuations in the atoms-cavity coupling on top of the mean-field treatment (see Sec. III.1). This method was developed and benchmarked against numerically exact results for bosonic atoms Bezvershenko et al. (2021); Halati et al. (2022); Halati (2021) and extended to fermionic atoms in Ref. Tolle et al. (2025a).
@@
with $\beta\!=\!1/k_{B}T$ the inverse temperature and $Z$ the partition function. In Ref. Bezvershenko et al. (2021) this approach was validated in a wide parameter regime by a comparison to results obtained with numerically exact results obtained with matrix product state methods for bosonic atomic. For the fermionic lattice model, when the cavity field has a finite occupation the effective Hamiltonian $\hat{H}_{\text{eff}}$, Eq. (6) is the ionic Hubbard model. This model was shown to obey Gaussian orthogonal ensemble (GOE) statistics in the presence of the staggered potential De Marco et al. (2022) associated with a chaotic dynamics. Since this level statistics De Marco et al. (2022) is similar to the case of interacting bosons Kolovsky et al. (2004); Kollath et al. (2010), we expect the thermal Ansatz for the fermions to lead to reliable results for the steady states above the self-organization transition and for higher dimensions both below and above the transition. In contrast, in one-dimension below the transition, the effective model is the Hubbard model, which is an integrable model. For this special case, more reliable results could be obtained using a generalized Gibbs Ansatz, taking into account the conserved quantities. However, due to the non-local nature of the conservation laws in many situations the dynamics of the Hubbard model is still well described neglecting these, such that we expect that typical quantities still show the correct features.
@@
is evaluated for the thermal state $\hat{\rho}_{\text{at}}$. The self-consistently determined temperature can stabilize to different values as we vary the parameters throughout the phase diagram, as we discuss in Sec. V. In particular this effect is a crucial ingredient for the many-body cooling mechanisms and the fluctuation-induced bistability Tolle et al. (2025a). In the normal phase fluctuations in the atoms-cavity coupling are present despite the vanishing steady state cavity field.
@@ III.2.1 Large temperature limit
This result agrees with the result of inverse effective temperature, $\beta$, for bosons Bezvershenko et al. (2021) and is consistent with results based on the semiclassical limit for non-interacting atoms Asbóth et al. (2005); Schütz and Morigi (2014); Piazza and Strack (2014b). This relation shows that by increasing the separation of energy scales between the cavity and the atoms, i.e. increasing $\hbar\delta$ and $\hbar\varGamma$, the atoms are driven into states with large effective temperatures.
@@ IV.1 Ionic Hubbard Hamiltonian
Here the expectation values are evaluated for the state one is interested in, so for example the ground state, or the finite temperature state. For the ground state one can use $\langle\hat{\gamma}^{\dagger}_{\pm,k,\sigma}\hat{\gamma}_{\pm,k,\sigma}\rangle\!=\!1$ for the filled momenta and $\langle\hat{\gamma}^{\dagger}_{\pm,k,\sigma}\hat{\gamma}_{\pm,k,\sigma}\rangle\!=\!0$ otherwise. We make use of these results to obtain analytical insights into the self-organization transition for atoms which have no direct interaction between themselves and and to obtain approximate results for strongly interacting atoms.
@@ IV.2 Ground state phase diagram and excitations
We discuss the phase diagram and low lying excitations of the ionic Hubbard model in one dimension. Since distinct phases arise at different fillings, we consider here two different fillings $n\!=\!N/L\!=\!\{1/2,1\}$.
@@ V Results
In this section, we present our results on the steady state phase diagram of the atoms-cavity coupled system. The transversely pumped coupled atoms-cavity open system exhibits a self-ordering transition Ritsch et al. (2013); Mivehvar et al. (2021), which has been studied widely for bosonic and fermionic atoms Asbóth et al. (2005); Larson et al. (2008b, a); Niedenzu et al. (2011); Piazza and Strack (2014b); Keeling et al. (2014); Chen et al. (2014). Our results investigate the self-organization transition for the considered system of interacting fermions coupled to the optical cavity. We analyze in detail the case of quarter filling (Sec. V.1), complementing our work presented in Ref. Tolle et al. (2025a). Whereas the numerical results shown are obtained at n=1/2n\!=\!1/2, our analytical findings can be generalized to lower fillings $n\!\leq\!1/2$. Additionally, we present and analyze results at half filling n=1n\!=\!1 (Sec. V.2). Fundamentally different behaviors are expected for the two fillings.
We begin the discussion (see Sec. V.1) by focusing on the regime around the self-ordering transition at $g\!\sim\!g_{\text{cr}}$ at quarter filling. We contrast the T=0T\!=\!0 MF results to the approach which includes fluctuations in the atoms-cavity coupling (Sec. V.1.1). Analytical approximations for the phase transition are derived in several limits. Then we emphasize features originating from the different cavity-cooling (Sec. V.1.2) and many-body cooling mechanisms (Sec. V.1.3), that are only possible due to the presence of fluctuations. At quarter filling, this includes the fluctuations-induced bistability (Sec. V.1.4), which has been the focus of Ref. Tolle et al. (2025a). In Sec. V.2.1, we turn to the bistability occurring at half filling. In this case the self-organization transition is coupled with the onset of a bistabilitym which is already visible for the T=0T\!=\!0 MF results.
We compare our results from solving of Eq. (5) and Eq. (15) with exact diagonalization (ED) for finite systems in one dimension to those obtained in the weak-tunneling perturbation theory in the thermodynamic limit, Eqs. (29). The latter is also applicable for higher dimensions of the atomic lattice.